Suppose that the time it takes for Andrew, Betty, or Carol to paint a garage working alone is A, B, or C hours respectively, where A,B,and C are positive integers. If they paint the garage together in 1 hour and 15 minutes, then what are the possible values of A,B, and C?
Answer by JulietG(1812)

15 minutes is what portion of an hour? 15/60 = 1/4. One of the painters must take 4 hours. That leaves a whole hour left. Any combination that can equal a whole works.
For instance, let's assume that Andrew takes 4 hours (for all cases since we need a quarter hour.)
If A = 1/4, and B = 0, then C = 1/1 (total of 1 1/4 hours)
If A = 1/4, and B = 1/1, then C = 0 (not everyone is in agreement that 0 is an integer)
If A = 1/4, and B = 2, then C = 2 (1/4 + 1/2 + 1/2 = 1 1/4)
Are there any other WHOLE numbers? I didn't find any. Of course, any of them can be the 4 hours...